A fast-food restaurant has identified three primary groups willing to purchase its meals. However, customers are willing to purchase only one meal each. The table shows the total number of meals bought at three different prices. The number of consumers is the total number of buyers at each price level.

Group Number of consumers Willingness to pay

High 350 $5.50

Medium 500 $4

Low 750 $2

The restaurant can produce a meal with no fixed costs and a constant marginal cost of $1 per unit.

Part 1 If the restaurant charges only one price for the meal, it will charge $____ and receive a profit of $_____ .

Part 2 Assume the firm can distinguish between customers and has the ability to charge different customers different prices. If the restaurant can perfectly price-discriminate, the highest price it will charge is $________. The total profits under perfect price discrimination are $_______ .

Part 3 Assume the firm can distinguish between customers and has the ability to charge different customers different prices. If the restaurant charges two prices, it will charge $______ and receive a profit of $_______ .
Answer #1

Let us assume form charges $ 5.50 then profit = 350(5.50-1) = $ 1575

Now if firm charges $ 4, then profit =(350+500)*(4-1)= $ 2550

If the firm charges $ 2, then profit= (350+500+750)*(2-1) = $1700

Hence in order to maximize profit firm must charge $ 4 per unit .

Part 1: Price $ 4 per unit

profit = $ 2,550

Part 2. Highest price charged by firm = $ 5.50

Total profit = (5.5-1)*350 + (4-1)*500+(2-1)*750 = $ 3825

Part 3. The firm must charge $ 4 and 2

Profit = $ 3300

Please contact before rating up will be obliged to you for your generous support thank you.